# Dinomaly: The Less Is More Philosophy in Multi-Class Unsupervised Anomaly Detection

### v2

@@ 3.2 Comparison to Multi-Class UAD SoTAs
We compare the proposed Dinomaly with the most advanced methods. Among them, RD4AD [2] based on feature reconstruction, SimpleNet [13] based on feature-level pseudo-anomaly, and DeSTSeg [12] based on feature reconstruction & pseudo anomaly are designed for conventional class-separated UAD settings. UniAD based on feature reconstruction, ReContrast [18] based on contrastive reconstruction, ViTAD [17] based on feature reconstruction & Transformer, DiAD [49] based on Diffusion reconstruction, and MambaAD [19] based on feature reconstruction & Mamba are designed for MUAD settings. Notably, ViTAD and MambaAD are contemporary arxiv preprints released within months. The intuitive comparison is already presented in Figure 1.
Experimental results are presented in Table 1, where Dinomaly surpasses compared methods by a large margin on all datasets and all metrics. On the most widely used MVTec-AD, Dinomaly produces image-level performance of 99.6/99.8/99.0 and pixel-level performance of 98.4/69.3/69.2/94.8, outperforming previous SoTAs by 1.0/0.2/1.2 and 0.7/9.1/7.7/1.6. This result declares that the image-level performance on the MVTec-AD dataset is nearly saturated under the MUAD setting. On the popular VisA, Dinomaly achieves image-level performance of 98.7/98.9/96.2 and pixel-level performance of 98.7/53.2/55.7/94.5, outperforming previous SoTAs by 3.2/2.5/4.2 and 0.2/5.3/5.1/2.6. On the Real-IAD that contains 30 classes, each with 5 camera views, we produce image-level and pixel-level performance of 89.3/86.8/80.2 and 98.8/42.8/47.1/93.9, outperforming previous SoTAs by 3.0/2.2/3.2 and 0.3/4.9/5.4/3.4, indicating our scalability to extremely complex scenarios. Per-class performances and qualitative visualization are presented in Appendix A.5 and A.6. In addition, adopting a larger backbone further improves the above performances, as presented in Table A2.
3.3 Comparison to Class-Separated UAD SoTAs
We also compare our Dinomaly with class-separated SoTAs, as shown in Table 2. On MVTec-AD and VisA, our Dinomaly under MUAD setting is comparable to conventional SoTAs that build individual models for each class [2; 13; 8; 15]. In addition, Dinomaly is subjected to nearly no performance drop compared to its class-separated counterpart on these datasets. On the complicated Real-IAD that involves more images, classes, and views, class-separated Dinomaly sets new SoTA records. Multi-class Dinomaly presents moderate performance drop but is still comparable to class-separated SoTAs.
@@ 3.3 Comparison to Class-Separated UAD SoTAs
I-AUROC P-AP P-AUPRO I-AUROC P-AP P-AUPRO I-AUROC P-AP P-AUPRO
Dinomaly (MUAD) 99.6 69.3 94.8 98.7 53.2 94.5 89.3 42.8 93.9
Dinomaly 99.7 68.9 95.0 98.9 50.7 95.1 92.0 45.2 95.1
RD4AD [2] 98.5 58.0 93.9 96.0 27.7 70.9 87.1 n/a 93.8
PatchCore [8] 99.1 56.1 93.5 95.1 40.1 91.2 89.4 n/a 91.5
SimpleNet [13] 99.6 54.8 90.0 96.8 36.3 88.7 88.5 n/a 84.6
EfficientAD [15] 99.1 63.8 93.5 98.1 40.8 94.0 n/a n/a n/a
Table 3: Ablations of Dinomaly elements on MVTec-AD (%). NB: Noisy Bottleneck. LA: Linear Attention. LC: Loose Constraint (2 groups). LL: Loose Loss. Results on VisA see Table A1.
@@ 3.4 Ablation Study
Overall Ablation. We conduct experiments to verify the effectiveness of the proposed elements, i.e., Noisy Bottleneck (NB), Linear Attention (LA), Loose Constraint (LC), and Loose Loss (LL). The already-powerful baseline is Dinomaly with noiseless MLP bottleneck, Softmax Attention, dense layer-to-layer supervision, and global cosine loss [18]. Results on MVTec-AD and VisA are shown in Table 3 and Table A1, respectively. NB and LL can directly contribute to the model performance. LA and LC boost the performance with the presence of NB. The use of LC is not solely beneficial because LC makes the reconstruction too easy without injected noise. Combining some of the proposed elements boosts the performance of the baseline, while employing them all produces the best results. Noisy Rates. We conduct ablations on the discarding rate of the Dropouts in MLP bottleneck, as shown in Table 4. Experimental results demonstrate that Dinomaly is robust to different levels of dropout rate. Reconstruction Constraint. We quantitatively examine different reconstruction schemes presented in Figure 4. As shown in Table 5, group-to-group LC outperforms layer-to-layer supervision. On image-level metrics, 1-group LC with all layers added performs similarly to its 2-group counterpart that separates low-level and high-level layers; however, 1-group LC mixes low-level and high-level features which is harmful for anomaly localization. More ablations on scalability, input size, pre-trained foundations, etc., are presented in Appendix A.3.
Table 4: Ablations of Dropout rates in Noisy Bottleneck, conducted on MVTec-AD (%). †: default.
Dropout rate Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
0 (noiseless) 98.19 99.55 98.51 97.55 63.11 64.39 93.33
0.1 99.54 99.75 98.90 98.35 69.46 69.19 94.53
0.2 † 99.60 99.78 99.04 98.35 69.29 69.17 94.79
0.3 99.65 99.83 99.16 98.34 68.46 68.81 94.63
0.4 99.64 99.80 99.23 98.22 67.95 68.33 94.57
0.5 99.56 99.81 99.14 98.15 67.43 67.82 94.64
Table 5: Ablations of reconstruction constraint, conduected on MVTec-AD (%). †: default.
Constraints Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
group-to-group (1 group) 99.64 99.80 99.36 98.18 64.79 65.40 93.96
group-to-group (2 groups)† 99.60 99.78 99.04 98.35 69.29 69.17 94.79
@@ 4 Conclusion
Appendix A Appendix / supplemental material
A.1 Related Work
Epistemic methods are based on the assumption that the networks respond differently during inference between seen input and unseen input. Within this paradigm, pixel reconstruction methods assume that the networks trained on normal images can reconstruct anomaly-free regions well, but poorly for anomalous regions. Auto-encoder (AE) [50; 51], variational auto-encoder (VAE) [52; 53], or generative adversarial network (GAN) [54; 55] are used to restore normal pixels. However, pixel reconstruction models may also succeed in restoring unseen anomalous regions if they resemble normal regions in pixel values or the anomalies are barely noticeable [2]. Therefore, feature reconstruction is proposed to construct features of pre-trained encoders instead of raw pixels [2; 3; 4]. To prevent the whole network from converging to a trivial solution, the parameters of the encoders are frozen during training. In feature distillation [5; 6], the student network is trained from scratch to mimic the output features of the pre-trained teacher network with the same input of normal images, also based on the similar hypothesis that the student trained on normal samples only succeed in mimicking features of normal regions.
Feature statistics methods [7; 8; 56; 9] memorize all normal features (or their modeled distribution) extracted by networks pre-trained on large-scale datasets and match them with test samples during inference. Since these methods require memorizing, processing, and matching nearly all features from training samples, they are computationally expensive in both training and inference, especially when the training set is large.
Multi-Class UAD. UniAD [3] first introduced multi-class anomaly detection, aiming to detect anomalies for different classes using a unified model. In this setting, conventional UAD methods often face the challenge of "identical shortcuts", where both anomaly-free and anomaly samples can be effectively recovered during inference [3]. It is caused by the diversity of multi-class normal patterns that drive the network to generalize on unseen patterns. This contradicts the fundamental assumption of epistemic methods. Many current researches focus on addressing this challenge [3; 24; 18; 57; 25]. UniAD [3] employs a neighbor-masked attention module and a feature-jitter strategy to mitigate these shortcuts. HVQ-Trans [24] proposes a vector quantization (VQ) Transformer model that induces large feature discrepancies for anomalies. LafitE [25] utilizes a latent diffusion model and introduces a feature editing strategy to alleviate this issue. DiAD [26] also employs diffusion models to address multi-class UAD settings. OmniAL [23] focuses on anomaly localization in the unified setting, preventing identical reconstruction by using synthesized pseudo anomalies. ViTAD [58] abstracts a unified feature-reconstruction UAD framework and employ Transformer building blocks. MambaAD [19] explores the recently proposed State Space Model (SSM), Mamba, in the context of multi-class UAD.
A.2 Full Implementation Details
ViT-Base/14 (patch size=14) pre-trained by DINOv2 with registers (DINOv2-R) [39] is utilized as the encoder. The discard rate of Dropout in Noisy Bottleneck is 0.2 by default, which is increased to 0.4 for the diverse Real-IAD. Loose constraint with 2 groups and $\mathcal{L}_{global-hm}$ loss are used by default. The input image is first resized to $448^{2}$ and then center-cropped to $392^{2}$, so that the feature map ($28^{2}$) is large enough for localization. StableAdamW optimizer [46] with AMSGrad [47] is utilized with $lr$ (learning rate)=2e-3, $\beta$=(0.9,0.999), $wd$ (weight decay)=1e-4 and $eps$=1e-10. The network is trained for 10,000 iterations for MVTec-AD and VisA and 50,000 iterations for Real-IAD under MUAD setting. The network is trained for 5,000 iterations on each class under the class-separated UAD setting. The $lr$ warms up from 0 to 2e-3 in the first 100 iterations and cosine anneals to 2e-4 throughout the training. The discarding rate in Equation 5 linearly rises from 0% to 90% in the first 1,000 iterations as warm-up (500 iters for class-separated setting). The anomaly map is obtained by upsampling the point-wise cosine distance between encoder and decoder feature maps (averaging if more than one pair or group). The mean of the top 1% pixels in an anomaly map is used as the image anomaly score. All experiments are conducted with random seed=1 with cuda deterministic for invariable weight initialization and batch order. Codes are implemented with Python 3.8 and PyTorch 1.12.0 cuda 11.3, and run on NVIDIA GeForce RTX3090 GPUs (24GB).
A.3 Additional Ablation Studies and Experiments
NB LA LC LL Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
95.81 96.35 92.06 97.97 47.88 52.55 93.43
97.38 97.74 94.07 97.84 50.42 54.57 93.71
95.74 96.23 91.87 98.01 47.89 52.58 93.34
96.39 97.01 92.54 97.37 46.80 51.66 92.75
96.93 97.26 93.32 98.37 49.52 53.59 94.11
97.52 97.75 94.33 98.06 51.49 55.09 93.75
98.06 98.37 95.18 98.21 51.43 54.89 93.94
98.57 98.77 95.75 98.57 52.29 55.38 94.28
98.22 98.43 95.27 98.51 53.11 55.48 94.24
98.73 98.87 96.18 98.74 53.23 55.69 94.50
Scalability. Previous works [3; 2; 17] reported that AD methods do not follow the model "scaling law", i.e., larger models do not necessarily produce better performance. For example, RD4AD [2] found WideResNet50 better than WideResNet101 as the encoder backbone. ViTAD [17] found ViT-Small better than ViT-Base. We conduct experiments to probe the influence of the scale of backbone Transformers in Dinomaly. ViT-Small, ViT-Base (default), and ViT-Large pre-trained by DINOv2-R are used as the encoder, respectively. ViT-Small has 12 layers, so we take the [3,4,5,…10]th layer as the interested 8 middle layers, which is the same as default ViT-Base. ViT-Large has 24 layers, so we take the [5,7,9,…19]th layer as the interested 8 middle layers. The layer hyperparameters of the decoder, such as embedding dimension and numbers of attention heads, follow the hyperparameters of the corresponding encoder. Other training strategies are identical to default. As shown in Table A2, the MUAD performance of Dinomaly follows the "scaling law". Dinomaly equipped with ViT-Small already produces state-of-the-art results. ViT-Large further boosts Dinomaly to an unprecedented higher record.
Table A2: Comparison of different ViT architectures, conducted on MVTec-AD (%). Latency per image is measured on NVIDIA RTX3090 with batch size=16.
Arch. Parameters MACs Latency Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
ViT-Small 37.4M 26.3G 6.8ms 99.26 99.67 98.72 98.07 68.29 67.78 94.36
ViT-Base 148.0M 104.7G 17.2ms 99.60 99.78 99.04 98.35 69.29 69.17 94.79
ViT-Large 275.3M 413.5G 41.3ms 99.77 99.92 99.45 98.54 70.53 70.04 95.09
@@ A.3 Additional Ablation Studies and Experiments
Image Size MACs Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
[…]
Pre-Trained Foundations. The representation quality of the frozen backbone Transformer is of great significance to unsupervised anomaly detection. We conduct extensive experiments to probe the impact of different pre-training methods, including supervised learning and self-supervised learning. DeiT [59] is trained on ImageNet[1] in a supervised manner by distilling CNNs. MAE [33], BEiTv2 [35], and D-iGPT [60] are based on masked image modeling (MIM). Given input images with masked patches, MAE [33] is optimized to restore raw pixels; BEiTv2 [35] is trained to predict the token index of VQ-GAN and CLIP; D-iGPT [60] is trained to predict the features of CLIP model. DINO [29] is based on positive-pair contrastive learning (CL), which is also referred to as self-distillation. It trains the network to produce similar feature representations given two views (augmentations) of the same image. iBot [36] and DINOv2 [30] combine MIM and CL strategies, marking the SoTA of self-supervised foundation models. DINOv2-R [39] is a variation of DINOv2 that employs 4 extra register tokens.
Table A4: Comparison between pre-trained ViT foundations, conducted on MVTec-AD (%). All models are ViT-Base. The patch size of DINOv2 and DINOv2-R is $14^{2}$; others are $16^{2}$.
Pre-Train Method Type Image Size Image-level Pixel-level

### v5

@@ 3.1 Dinomaly Framework
[…]
3.4 Loose Reconstruction
“The tighter you squeeze, the less you have.”
Loose Constraint. Pioneers of feature-reconstruction/distillation UAD methods [46, 10] are inspired by knowledge distillation [20]. Most reconstruction-based methods distill specific encoder layers (e.g. 3 last layers of 3 ResNet stages) by the corresponding decoder layers [10, 46, 65] (Figure 4(a)) or the last decoder layer [60, 58] (Figure 4(b)). Intuitively, with more encoder-decoder feature pairs (Figure 4(c)), UAD model can utilize more information in different layers to discriminate anomalies. However, according to the intuition of knowledge distillation, the student (decoder) can better mimic the behavior of the teacher (encoder) given more layer-to-layer supervision [30], which is harmful for UAD models that detect anomalies by encoder-decoder discrepancy. This phenomenon is also embodied as identity mapping. Thanks to the top-to-bottom consistency of columnar Transformer layers, we propose to loosen the layer-to-layer constraint by adding up all feature maps of interested layers as a whole group, as shown in Figure 4(d). This scheme can be seen as loosening the layer-to-layer correspondence and providing the decoder with more degrees of freedom, so that the decoder is allowed to act much more differently from the encoder when the input pattern is unseen. Because features of shallow layers contain low-level visual characters that are helpful for precise localization, we can further group the features into the low-semantic-level group and high-semantic-level group, as shown in Figure 4(e).
$\mathcal{L}_{global-hm}=\mathcal{D}_{cos}(\mathcal{F}(f_{E}),\mathcal{F}(\hat{f_{D}})),$ (4)
$\small\hat{f_{D}}(h,w)=\left\{\begin{array}[]{l}sg(f_{D}(h,w))_{0.1},\text{ if}\mathcal{D}_{cos}(f_{D},f_{E})<k\%_{batch}\\ f_{D}(h,w),\text{ else}\\ \end{array}\right.$ (5)
$\small\mathcal{D}_{cos}(a,b)=1-\frac{{a}^{T}\cdot b}{\left\|{a}\right\|~{}\left\|{b}\right\|},$ (6)
where $\mathcal{D}_{cos}$ denotes cosine distance, $\mathcal{F}(\cdot)$ denotes flatten operation, $f_{D}(h,w)$ represents the feature point at $(h,w)$, and $sg(\cdot)_{0.1}$ denotes shrink the gradient to one-tenth of the original  [^2] Complete stop-gradient causes optimization instability occasionally.. $\mathcal{D}_{cos}(f_{D}(h,w),f_{E}(h,w))<k\%_{batch}$ selects $k\%$ feature points with smaller cosine distance within a batch for gradient shrinking. Total loss is the average $\mathcal{L}_{global-hm}$ of all encoder-decoder feature pairs.
4 Experiments
4.1 Experimental Settings
Datasets. MVTec-AD [3] contains 15 objects (5 texture classes and 10 object classes) with a total of 3,629 normal images as the training set and 1,725 images as the test set (467 normal, 1,258 anomalous). VisA [70] contains 12 objects. Training and test sets are split following the official splitting, resulting in 8,659 normal images in the training set and 2,162 images in the test set (962 normal, 1,200 anomalous). Real-IAD [54] is a large UAD dataset recently released, containing 30 distinct objects. We follow the official splitting that includes all views, resulting in 36,465 normal images in the training set and 114,585 images in the test set (63,256 normal, 51,329 anomalous).
Metrics. Following prior works [17, 65], we adopt 7 evaluation metrics. Image-level anomaly detection performance is measured by the Area Under the Receiver Operator Curve (AUROC), Average Precision (AP), and $F_{1}$ score under optimal threshold ($F_{1}$-max). Pixel-level anomaly localization is measured by AUROC, AP, $F_{1}$-max and the Area Under the Per-Region-Overlap (AUPRO). The results of a dataset is the average of all classes.
@@ 4.1 Experimental Settings
I-AUROC P-AUROC P-AUPRO I-AUROC P-AUROC P-AUPRO I-AUROC P-AUROC P-AUPRO
Dinomaly (MUAD) 99.6 98.4 94.8 98.7 98.7 94.5 89.3 98.8 93.9
Dinomaly 99.7 99.9 95.0 98.9 98.9 95.1 92.0 99.1 95.1
RD4AD [10] 98.5 97.8 93.9 96.0 90.1 70.9 87.1 n/a 93.8
PatchCore [45] 99.1 98.1 93.5 94.7 98.5 91.8 89.4 n/a 91.5
SimpleNet [34] 99.6 98.1 90.0 97.1 98.2 90.7 88.5 n/a 84.6
Table 3: Ablations of Dinomaly elements on MVTec-AD (%). NB: Noisy Bottleneck. LA: Linear Attention. LC: Loose Constraint (2 groups). LL: Loose Loss. As MVTec-AD has reached saturation, we also present the results on VisA (Table A4).
@@
Implementation Details. ViT-Base/14 (patchsize=14) pre-trained by DINOv2-R [7] is used as the encoder by default. The drop rate of Noisy Bottleneck is 0.2 by default and increases to 0.4 on the diverse Real-IAD. Loose constraint with 2 groups is employed, and the anomaly map is given by the mean per-point cosine distance of the 2 groups. The input image is first resized to $448^{2}$ and then center-cropped to $392^{2}$, so the feature map ($28^{2}$) is large enough for anomaly localization. StableAdamW optimizer [56] with AMSGrad [41] (more stable than AdamW [35] in training) is utilized with $lr$=2e-3, $\beta$=(0.9,0.999) and $wd$=1e-4. The network is trained for 10,000 iterations (steps) on MVTec-AD and VisA, and 50,000 iterations on Real-IAD. Detailed settings are available in Appendix B.
4.2 Comparison to Multi-Class UAD SoTAs
We compare the proposed Dinomaly with the most advanced UAD and MUAD methods [10, 34, 67, 14, 65, 18, 17]. Experimental results are presented in Table 1, where Dinomaly surpasses compared methods by a large margin on all datasets and all metrics. On the most widely used MVTec-AD, Dinomaly produces image-level performance of 99.6/99.8/99.0 (%) and pixel-level performance of 98.4/69.3/69.2/94.8, outperforming previous SoTAs by 1.0/0.2/1.2 and 0.7/9.1/7.7/1.6. This result declares that the image-level performance on the MVTec-AD dataset is nearly saturated under the MUAD setting. On the popular VisA, Dinomaly achieves image-level performance of 98.7/98.9/96.2 and pixel-level performance of 98.7/53.2/55.7/94.5, outperforming previous SoTAs by 3.2/2.5/4.2 and 0.2/5.3/5.1/2.6. On the Real-IAD that contains 30 classes, each with 5 camera views, we produce image-level and pixel-level performance of 89.3/86.8/80.2 and 98.8/42.8/47.1/93.9, outperforming previous SoTAs by 3.0/2.2/3.2 and 0.3/4.9/5.4/3.4, indicating our scalability to extremely complex scenarios. Per-class performances and qualitative visualization are presented in Appendix E and F. We also produce superior results on other popular UAD benchmarks, i.e., MPDD [24], BTAD [38], and Uni-Medical [66], with I-AUROC of 97.2, 95.4, and 84.9, respectively, as shown in Table A13 in Appendix.
4.3 Comparison to Class-Separated UAD SoTAs
Dinomaly is also compared with class-separated SoTAs, as shown in Table 2. Dinomaly under MUAD setting is comparable to conventional methods [10, 34, 45] that build individual models for each class. On MVTec-AD and VisA, multi-class Dinomaly (first row) is subjected to nearly no performance drop compared to its class-separated counterpart (second row). On the complicated Real-IAD that involves more classes and views, multi-class Dinomaly suffers a moderate performance drop but is still comparable to class-separated SoTAs.
Table 4: Scaling of ViT model sizes on MVTec-AD (%). Im/s (Troughoutput, image per second) is measured on NVIDIA RTX3090 with batch size=16. Results on VisA and Real-IAD are shown in Table A3. †:default.
Arch. Params MACs Im/s Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
ViT-Small 37.4M 26.3G 153.6 99.26 99.67 98.72 98.07 68.29 67.78 94.36
ViT-Base† 148.0M 104.7G 58.1 99.60 99.78 99.04 98.35 69.29 69.17 94.79
ViT-Large 275.3M 413.5G 24.2 99.77 99.92 99.45 98.54 70.53 70.04 95.09
Table 5: Scaling input size on MVTec-AD (%). †: default. Compared methods yield degradation when increasing input size.
Method Input Size Image-Level Pixel-Level
RD4AD $256^{2}$ 94.6/96.5/96.1 96.1/48.6/53.8/91.1
$320^{2}$ 93.2/96.9/95.6 95.7/55.1/57.5/91.1
$384^{2}$ 91.9/96.2/95.0 94.9/52.1/55.3/90.8
ReContrast $256^{2}$ 98.3/99.4/97.6 97.1/60.2/61.5/93.2
$320^{2}$ 98.2/99.2/97.5 96.8/61.8/62.6/93.3
$384^{2}$ 95.2/98.0/96.4 96.5/57.7/59.5/92.6
Dinomaly $280^{2}$ 99.6/99.8/99.3 98.2/65.2/66.3/93.6
$336^{2}$ 99.6/99.8/99.2 98.3/67.2/67.8/94.2
$392^{2}$ 99.6/99.8/99.0 98.4/69.3/69.2/94.8
@@ 4.4 Ablation Study
Overall Ablation. We conduct experiments to verify the effectiveness of the proposed elements, i.e., Noisy Bottleneck (NB), Linear Attention (LA), Loose Constraint (LC), and Loose Loss (LL). The already-powerful baseline (first row) is Dinomaly with noiseless MLP bottleneck, Softmax Attention, dense layer-to-layer supervision, and global cosine loss. This baseline is very similar to ViTAD [64] and the ViT version of RD4AD [10]. Results on MVTec-AD and VisA are shown in Table 3 and Table A4, respectively. NB and LL can directly contribute to the model performance. LA and LC boost the performance with the presence of NB. The use of LC is not solely beneficial because LC makes the reconstruction too easy without injected noise.
[…]
Input Scalability. Though it seems unfair to compare Dinomaly with previous works that take smaller images as input, we contend that increasing their input size not only fails to benefit but actively undermines their performance, especially for image-level detection performance, as shown in Table 5. Therefore, we follow the common comparison strategy based on ”optimum vs. optimum”. On the contrary, Dinomaly enjoys scaling input size for anomaly localization, while still producing SoTA performance given smaller images. Details are presented in Table A2 in Appendix.
Figure 5: Image-level AUROC of Dinomaly equipped with various ViT foundations, and their linear-probing accuracy on ImageNet. MIM: Masked Image Modeling. CL: Contrastive Learning.
ViT Foundations. We conduct extensive experiments to investigate the impact of diverse pre-trained ViT foundations, including DeiT [50], MAE [19], D-iGPT [44], MOCOv3 [6], DINO [4], iBot [69], DINOv2 [39], and DINOv2-R [7]. As shown in Figure 5, Dinomaly is robust to the choice of backbone. Almost all foundation models can produce SoTA-level results with image-level AUROC higher than 98%. The only notable exception is MAE, which, without fine-tuning, was reported to be less effective across various unsupervised tasks, e.g, kNN and linear-probing [39]. The optimal input size varies because the these backbones are pre-trained on different resolutions. Interestingly, we found the anomaly detection performance to be strongly correlated with the accuracy of ImageNet linear-probing (freeze backbone & only tune linear classifier) of the foundation model, suggesting the possibility of further improvement by simply adopting a future foundation model. Detailed results and analysis are presented in Appendix and Table A1.
Additional experiments and results are detailed in the Appendix C, encompassing evaluations of various pre-trained foundations, ablation studies of each components, hyperparameter optimization, and other in-depth analyses.
@@ 5 Conclusion
Acknowledgments
The authors acknowledge supports from National Natural Science Foundation of China (U22A2051, 82027807), National Key Research and Development Program of China (2022YFC2405200), Tsinghua-Foshan Innovation Special Fund (2021THFS0104), and Institute for Intelligent Healthcare, Tsinghua University (2022ZLB001).
Supplementary Material
@@ Acknowledgments
Pre-Train Backbone Type Image Size Image-level Pixel-level
@@
[…]
Image Size MACs Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
[…]
A Additional Related Work
Here, we discussed general methods for unsupervised anomlay detection. Epistemic methods are based on the assumption that the networks respond differently during inference between seen input and unseen input. Within this paradigm, pixel reconstruction methods assume that the networks trained on normal images can reconstruct anomaly-free regions well, but poorly for anomalous regions. Auto-encoder (AE) [2, 63], variational auto-encoder (VAE) [32, 9], or generative adversarial network (GAN) [47, 1] are used to restore normal pixels. However, pixel reconstruction models may also succeed in restoring unseen anomalous regions if they resemble normal regions in pixel values or the anomalies are barely noticeable [10]. Therefore, feature reconstruction is proposed to construct features of pre-trained encoders instead of raw pixels [10, 60, 58]. To prevent the whole network from converging to a trivial solution, the parameters of the encoders are frozen during training. In feature distillation [46, 55], the student network is trained from scratch to mimic the output features of the pre-trained teacher network with the same input of normal images, also based on the similar hypothesis that the student trained on normal samples only succeed in mimicking features of normal regions.
Feature statistics methods [8, 45, 49, 27] memorize all normal features (or their modeled distribution) extracted by networks pre-trained on large-scale datasets and match them with test samples during inference. Since these methods require memorizing, processing, and matching nearly all features from training samples, they are computationally expensive in both training and inference, especially when the training set is large.
B Full Implementation Details
ViT-Base/14 (patch size=14) pre-trained by DINOv2 with registers (DINOv2-R) [7] is utilized as the encoder by default. The discard rate of Dropout in Noisy Bottleneck is 0.2 by default, which is increased to 0.4 for the diverse Real-IAD. Loose constraint with 2 groups and $\mathcal{L}_{global-hm}$ loss are used by default. The input image is first resized to $448^{2}$ and then center-cropped to $392^{2}$, so that the feature map ($28^{2}$) is large enough for localization. As previously discussed, the middle 8 layers of 12-layer ViT-Base are used for reconstruction and feeding the bottleneck. ViT-Small also has 12 layers, which is the same. ViT-Large contains 24 layers; therefore, we use the [4,6,8,…18] layers (index start from 0). The decoder always contains 8 layer.
StableAdamW optimizer [56] with AMSGrad [41] is utilized with $lr$ (learning rate)=2e-3, $\beta$=(0.9,0.999), $wd$ (weight decay)=1e-4 and $eps$=1e-10. The network is trained for 10,000 iterations for MVTec-AD and VisA and 50,000 iterations for Real-IAD under MUAD setting. The network is trained for 5,000 iterations on each class under the class-separated UAD setting. The $lr$ warms up from 0 to 2e-3 in the first 100 iterations and cosine anneals to 2e-4 throughout the training. The discarding rate in Equation 5 linearly rises from 0% to 90% in the first 1,000 iterations as warm-up (500 iters for class-separated setting). The anomaly map is obtained by upsampling the point-wise cosine distance between encoder and decoder feature maps (averaging if more than one pair or group). The mean of the top 1% pixels in an anomaly map is used as the image anomaly score. All experiments are conducted with random seed=1 with cuda deterministic for invariable weight initialization and batch order. Codes are implemented with Python 3.8 and PyTorch 1.12.0 cuda 11.3, and run on NVIDIA GeForce RTX3090 GPUs (24GB).
Most results of compared MUAD SoTAs are directly drawn from a benchmark paper ADer [66]. We express great thanks for their wonderful work.
Table A3: Scaling of ViT architectures on VisA and Real-IAD (%). †: default.
Dateset Arch. Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
VisA [70] ViT-Small 97.94 98.09 95.33 98.57 51.19 55.10 93.71
ViT-Base† 98.73 98.87 96.18 98.74 53.23 55.69 94.50
ViT-Large 98.85 99.09 96.12 99.10 55.68 57.33 94.76
Real-IAD [54] ViT-Small 89.10 86.91 79.87 98.69 41.88 46.74 94.08
ViT-Base† 89.33 86.77 80.17 98.84 42.79 47.10 93.86
ViT-Large 90.07 87.57 80.90 99.02 44.29 48.36 94.37
NB LA LC LL Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUPRO
95.81 96.35 92.06 97.97 47.88 52.55 93.43
97.38 97.74 94.07 97.84 50.42 54.57 93.71
95.74 96.23 91.87 98.01 47.89 52.58 93.34
96.39 97.01 92.54 97.37 46.80 51.66 92.75
96.93 97.26 93.32 98.37 49.52 53.59 94.11
97.52 97.75 94.33 98.06 51.49 55.09 93.75
98.06 98.37 95.18 98.21 51.43 54.89 93.94
98.57 98.77 95.75 98.57 52.29 55.38 94.28
98.22 98.43 95.27 98.51 53.11 55.48 94.24
98.73 98.87 96.18 98.74 53.23 55.69 94.50
C Additional Ablation and Experiment
Pre-Trained Foundations. The representation quality of the frozen backbone Transformer is of great significance to unsupervised anomaly detection. We conduct extensive experiments to probe the impact of different pre-training methods, including supervised learning and self-supervised learning. DeiT [50] is trained on ImageNet[11] in a supervised manner by distilling CNNs. MAE [19], BEiTv2 [40], and D-iGPT [44] are based on masked image modeling (MIM). Given input images with masked patches, MAE [19] is optimized to restore raw pixels; BEiTv2 [40] is trained to predict the token index of VQ-GAN and CLIP; D-iGPT [44] is trained to predict the features of CLIP model. MOCOv3 [6] is based on contrastive learning (CL), pulling the representations of the similar images and pushing those of different images. DINO [4] is based on positive-pair contrastive learning, which is also referred to as self-distillation. It trains the network to produce similar feature representations given two views (augmentations) of the same image. iBot [69] and DINOv2 [39] combine MIM and CL strategies, marking the SoTA of self-supervised foundation models. DINOv2-R [7] is a variation of DINOv2 that employs 4 extra register tokens.